Construct a regular expression that recognizes the following language of strings over the alphabet {0 1}:
The language consisting of the set of all bit strings that contain two or three symbols.
Answer #1

answer)

given alphabet is {0,1}

set of all bit strings with two symbols is constructed by (0+1)(0+1)=00+01+10+11

set of all bit strings with three symbols is constructed by (0+1)(0+1)(0+1)=000+001+010+011+100+101+110+111

regular expression for set of all bit string that contain two or three symbols is union of above two regular expressions.

i.e (0+1)(0+1)+(0+1)(0+1)(0+1)